OpenStudy (anonymous):

Which of the following is the correct factored form of the given equation? x^2+3x-4=0 A. (x - 1)(x + 4) = 0 B. (x - 1)(x - 4) = 0 C. (x + 1)(x - 4) = 0

OpenStudy (anonymous):

@nicholeykeysj

OpenStudy (anonymous):

A

OpenStudy (anonymous):

i dont want direct answers

OpenStudy (anonymous):

b

OpenStudy (anonymous):

i want to know how to get the answer

OpenStudy (anonymous):

nichole

OpenStudy (anonymous):

B is wrong. A is correct

OpenStudy (anonymous):



OpenStudy (anonymous):

(x-1)(x+4)=(x*x)+(x*4)+(-1*x)+(-1*4) =x^2+4x-x-4 = x^2+3x-4

OpenStudy (anonymous):

@CountryGurl15

OpenStudy (anonymous):

I'm sorry i suck at factoring =)

OpenStudy (anonymous):

oh